# Equal contribution

### v1

@@ 5.1 Memorization of fine-tuning training data
After fine-tuning, 319 out of the 3,526 SD.Base examples are memorized. In Figure 2, we plot the (normalized) edit distance between the target and model completion on these 319 examples after RL fine-tuning, where we vary the $\alpha$ coefficient that determines the strength of KL regularization; a small $\alpha$ results in a small KL regularisation penalty, allowing the optimized policy to deviate away from its initialisation (FT.1), while a large $\alpha$ has the opposite behavior. All three RL fine-tuned models maintain roughly the same memorization rate – between 43-47% of the 319 examples remain memorized – while the remaining examples have completions with an edit distance larger than 0.1. We observe that the RL fine-tuned model with $\alpha=0.005$ produces responses with larger edit distances compared to the models optimized with larger $\alpha$ values. This suggests that a smaller $\alpha$ allows the optimized policy model more latitude to produce new completions to prompts that have been memorized by the initial fine-tuned model, thereby reducing the rate of memorization that persists from the initial fine-tuned model. However, as Figure 2 shows, the impact of $\alpha$ on decreasing the persistence of memorization from the fine-tuned initialisation is marginal. We show some examples from SD.Base that remain memorized in Appendix F.
@@ 5.2 Memorization of reward model training data through RL fine-tuning
We next analyze the potential for the RLFT model to memorize reward model training data via the reward model signal used during the RL fine-tuning stage. First, we check the reward model has fit its training data well by measuring the distribution of reward scores it assigns to SD.Links. In Figure 2, we plot the scores of both positively and negatively labelled examples from SD.Links under three reward models: a reward model that was not trained on SD (RM.1), a reward model trained solely on SD (referred to as RM.2), and a reward model that included SD and other Python datasets in its training dataset (referred to as RM.3). Clearly, the reward models that included SD.Links in their training data assign high scores to positive examples and low scores to negative examples, while the two distributions of scores for examples with ground truth positive and negative labels, respectively are very close to one another when scored by a reward model that did not train on SD.Links.
We then evaluate if the RLFT model is able to memorize (and then regurgitate) the reward model’s training data after RL optimization. We perform RL fine-tuning on prompts from SD (we refer to this model as RLFT.2), where we use RM.3 as the reward model (this model was trained on SD). We initialize RLFT.2 from a model that was not fine-tuned on SD (referred to as FT.2). This is to rule out the possibility that we see memorization of SD examples after RL fine-tuning because of memorization in the fine-tuning stage. We will contrast the memorization rates of RLFT.2 against a model that is directly fine-tuned on SD (referred to as FT.3). This gives us a comparison between memorization rates from fine-tuning directly on a dataset and from RL fine-tuning (where that dataset is used for reward model training). Note, although SD is also used for RL fine-tuning in this experiment (RLFT.2), only the prompts are used, while the associated targets are discarded.
@@
One may hypothesize that we see fairly little memorization of the reward model training examples from RL fine-tuning because the reward is a weak signal with which to propagate memorization during RL fine-tuning. However, this is not the only factor that could impact the difficulty of propagating memorization of the reward model’s training data to the RL fine-tuned model. The KL regularization penalty, which encourages the RL fine-tuned model to stay close to its initialization, also constrains the model from over-optimizing the reward model (and thereby decreases the chance that it memorizes reward model training data).
We measure how KL regularization impacts the memorization of reward model training data by sweeping over the KL regularization coefficient used in the RL fine-tuning step, $\alpha$. We use RLFT.2 for this experiment, which was optimized against RM.3 (a reward model that included SD in its training dataset). With a small KL penalty ($\alpha=0.005$), the RL fine-tuning procedure regresses to always output the string #. We hypothesize that the reason for this is as follows: the targets on SD.Base are selected by randomly choosing a line in the example, and Python comments in SD.Base examples account for  40-50% of lines. Additionally, all examples in SD.Base are positively labelled, so it’s very likely that the reward model heavily associates Python comments with a positive score.
@@ 5.3 Memorization from RL fine-tuning
Finally, we measure the rate of prompt memorization in RL fine-tuning. We perform RL fine-tuning on SD (referred to as RLFT.4, i.e. during RLHF the model is exposed to SD only via the prompts during RL fine-tuning), and measure memorization taking examples with at least 100 tokens in the prompt, splitting them into a prompt and target of 50 tokens each, and match generations against this target (we use this length to align with the analysis in Carlini et al. (2022)). The results are shown in Figure 4. We also plot this as a function of the KL regularization coefficient $\alpha$ used in RL fine-tuning (please see Appendix E for a sensitive analysis on learning rates rather than $\alpha$). A smaller $\alpha$ allows the RLFT model to overfit more to its training prompts resulting in increased memorization. However, even with $\alpha=0.005$, fewer than 0.5% of the RL training prompts are memorized after 70 epochs of RLFT.
@@ 6 Discussion
As our experimental results have shown, and previous work has supported (Carlini et al., 2022), large language models can memorize training data. However, modern deployed language models go through extensive post-training after pre-training, with most post-training primarily relying on RLHF to align with human preferences. To date, there has been little analysis of the impact of this post-training phase on memorization of training data. Our study performs a careful analysis of the potential for memorization of data introduced in each of the three phases of RLHF. While Rando and Tramèr (2024) examines adversarial poisoning of the RLHF pipeline, our study examines how memorization can arise and propagate among the stages of RLHF in the average case, which is likely more representative of how large model providers currently align their models with user preferences.
@@
It’s important to note that measuring the risk of sensitive data leakage requires more nuance than assessing verbatim memorization because it requires knowledge of exactly what data is considered sensitive. This becomes inherently more difficult with larger, real-world datasets that may contain sensitive information not known ahead of time, in contrast to our work where we create the sensitive data ourselves. Additionally, what defines memorization is dependent on the data domain. In the context of code completion, we found that measuring counterfactual memorization was critical, since the highly structured nature of code introduces many potential confounding explanations of why a model produces an output that matches a given target (e.g., the model is a good code completion model). We also found it important to exclude examples where the target was approximately contained in the prompt, since code contains information that is often repeated (e.g. variable names and method calls). Restricting the impact of these confounders on a memorization analysis requires careful consideration of the unique qualities of the data.
There are multiple directions of future work that remain interesting. We are curious to see whether post-training methods that replace RLHF with supervised learning directly on preference data, such as Direct Preference Optimization (DPO) (Rafailov et al., 2023), may exhibit increased memorization of reward model preference data, since methods like DPO fine-tune models directly on preference data as opposed to using it indirectly via the intermediary reward model as in RLHF. Additionally, since prior work suggests that the potential for memorization increases with scale (Carlini et al., 2022), it would be interesting to study whether larger models exhibit increased memorization of RM training data during RLFT, as we saw little evidence of this at the scale of our RLFT model (1.8 B).
@@ A.4 Benchmark evaluation
Additionally, we evaluate our model on the canonical Python coding evaluation of HumanEval introduced by  Chen et al. [2021] that measures correctness for creating programs from docstrings. While not a code completion specific dataset, oftentimes users find value in the ability of a model to generate a function given a function header, docstring, or commented description. This benchmark serves as a proxy for that task. Overall, most models we train are competitive with other small models on code completions tasks (see Appendix G).
@@ Appendix B Instantiation of the working definitions of memorization
Code is highly structured and a model that has learnt language specific style guidelines and has access to relevant contextual information from an examples prompt could conceivably output a completion similar to the target without having memorized the example during training. We control for this by utilizing the counterfactual definition of memorization in Definition 3.1. Comparing completions under the model we are inspecting for memorization and under a control model (that has a separate training dataset) will allow us to rule out matches that are not due to memorization, since the control model has a separate training dataset. We then only count an example as memorized if it is $k$-approximately counterfactually memorized using this fixed control model. Throughout all experiments in Section 5, we use FT.2 as a control model, as this model was not trained on data from SD but performs well on benchmark metrics. We refer to Appendix G for full details of performance and training data for FT.2.
@@ Appendix E Further experiments on RL prompt memorization
Appendix F Selection of memorized examples
@@ 12 Returns:
Appendix G Benchmark results and training data provenance summary
@@
Table 1: Training data provenance details for the models we train. All fine-tuned models are initialized from the Gemini Nano-1 (1.8B) pre-trained on 14B tokens from public code repositories as described in Section A.3. SD refers to the synthetic dataset we use for memorization analysis and is described in Appendix H, and PD refers to the public dataset we use to maintain model utility and is described in Section A.1.
@@
Appendix H Construction of the synthetic dataset (SD) and a selection of examples

### v2

@@ 5.1 Memorization of fine-tuning training data
After fine-tuning, 319 out of the 3,526 SD.Base examples are memorized. In Figure 2, we plot the (normalized) edit distance between the target and model completion on these 319 examples after RL fine-tuning, where we vary the $\alpha$ coefficient that determines the strength of KL regularization; a small $\alpha$ results in a small KL regularisation penalty, allowing the optimized policy to deviate away from its initialisation (FT.1), while a large $\alpha$ has the opposite behavior. All three RL fine-tuned models maintain roughly the same memorization rate – between 43-47% of the 319 examples remain memorized – while the remaining examples have completions with an edit distance larger than 0.1. We observe that the RL fine-tuned model with $\alpha=0.005$ produces responses with larger edit distances compared to the models optimized with larger $\alpha$ values. This suggests that a smaller $\alpha$ allows the optimized policy model more latitude to produce new completions to prompts that have been memorized by the initial fine-tuned model, thereby reducing the rate of memorization that persists from the initial fine-tuned model. However, as Figure 2 shows, the impact of $\alpha$ on decreasing the persistence of memorization from the fine-tuned initialisation is marginal. We show some examples from SD.Base that remain memorized in Appendix H.
@@ 5.2 Memorization of reward model training data through RL fine-tuning
We next analyze the potential for the RLFT model to memorize reward model training data via the reward model signal used during the RL fine-tuning stage. First, we check the reward model has fit its training data well by measuring the distribution of reward scores it assigns to SD.Links. In Figure 2, we plot the scores of both positively and negatively labelled examples from SD.Links under three reward models: a reward model that was not trained on SD (RM.1), a reward model trained solely on SD (referred to as RM.2), and a reward model that included SD and other Python datasets in its training dataset (referred to as RM.3). Clearly, the reward models that included SD.Links in their training data assign high scores to positive examples and low scores to negative examples, while the two distributions of scores for examples with ground truth positive and negative labels, respectively are very close to one another when scored by a reward model that did not train on SD.Links. We then evaluate if the RLFT model is able to memorize (and then regurgitate) the reward model’s training data after RL optimization. We perform RL fine-tuning on prompts from SD (we refer to this model as RLFT.2), where we use RM.3 as the reward model (this model was trained on SD). We initialize RLFT.2 from a model that was not fine-tuned on SD (referred to as FT.2). This is to rule out the possibility that we see memorization of SD examples after RL fine-tuning because of memorization in the fine-tuning stage. We will contrast the memorization rates of RLFT.2 against a model that is directly fine-tuned on SD (referred to as FT.3). This gives us a comparison between memorization rates from fine-tuning directly on a dataset and from RL fine-tuning (where that dataset is used for reward model training). Note, although SD is also used for RL fine-tuning in this experiment (RLFT.2), only the prompts are used, while the associated targets are discarded.
@@
One may hypothesize that we see little memorization of the reward model training examples from RL fine-tuning because the reward is a weak signal with which to propagate memorization during RL fine-tuning. However, this is not the only factor that could impact the difficulty of propagating memorization of the reward model’s training data to the RL fine-tuned model. The KL regularization penalty, which encourages the RL fine-tuned model to stay close to its initialization, also constrains the model from over-optimizing the reward model ( thereby decreases the chance that it memorizes reward model training data). We measure how KL regularization impacts the memorization of reward model training data by sweeping over the KL regularization coefficient used in the RL fine-tuning step, $\alpha$. We use RLFT.2 for this experiment, which was optimized against RM.3 (which included SD in its training dataset). With a small KL penalty ($\alpha=0.005$), the RL fine-tuning procedure regresses to always output the string #. We hypothesize that the reason for this is as follows: the targets on SD.Base are selected by randomly choosing a line in the example, and Python comments in SD.Base examples account for  40-50% of lines. Additionally, all examples in SD.Base are positively labelled, so it’s likely that the reward model heavily associates Python comments with a positive score.
@@ 5.3 Memorization from RL fine-tuning
We then measure the rate of prompt memorization in RL fine-tuning, the final stage of RLHF. We perform RL fine-tuning on SD (referred to as RLFT.4, i.e. during RLHF the model is exposed to SD only via the prompts during RL fine-tuning), and measure memorization taking examples with at least 100 tokens in the prompt, splitting them into a prompt and target of 50 tokens each, and match generations against this target (we use this length to align with the analysis in Carlini et al. (2022)). The results are shown in Figure 4. We also plot this as a function of the KL regularization coefficient $\alpha$ used in RL fine-tuning (please see Appendix E for a sensitive analysis on learning rates rather than $\alpha$). A smaller $\alpha$ allows the RLFT model to overfit more to its training prompts resulting in increased memorization. However, even with $\alpha=0.005$, fewer than 0.5% of the RL training prompts are memorized after 70 epochs of RLFT.
5.4 Memorization from direct preference learning via IPO
We next analyze the potential for training data memorization when directly learning from preference data via the IPO algorithm (Azar et al., 2024). We separately fine-tune Gemma 2B and 7B (Team et al., 2024) on SD.Links and SD.Base, and measure the counterfactual memorization rates compared to the base Gemma 2B and 7B models, respectively. Full results are given in Appendix F.
Results on SD.Base We find that after IPO, the Gemma 2B model memorizes 18.2% of the SD.Base and the Gemma 7B model memorizes 19.5% of SD.Base. In contrast, RLFT only memorizes 0.9% of these examples when they are included in reward model training, as discussed in section Section 5.2.
Results on SD.Links After IPO, both Gemma 2B and Gemma 7B memorize only 0.6% of SD.Links. In contrast, RLFT memorizes 0% of these examples when they are included in reward model training.
Effect of scale We see slightly increased memorization of SD.Base by the Gemma 7B model trained via IPO. This is consistent with previous literature that has suggested larger models have more capacity to memorize (Carlini et al., 2022). The effect of scaling the reward model and RLFT model during RLHF is discussed in section Section 5.5.
Together, these results strongly suggest that direct preference learning algorithms such as IPO, though described as an equivalent optimization objective to RLHF in the RLHF literature, exhibit stronger memorization of preference data than RLHF.
5.5 Results on different model sizes and tasks
Our initial focus on evaluating memorization on code completion tasks was motivated by the prevalence of code completion as a daily use case for large language models and by the potential for memorization to significantly impact model quality and training data privacy. However, we also investigate the generalizability of our findings across different model scales, task domains, and datasets, and find our conclusions to be robust. We increase model scales, by scaling the reward model from T5-Base (220M) to Gemini Nano-1 (1.8B) and the RLFT model from Gemini Nano-1 to Gemini Pro. We evaluate memorization on a different code completion dataset, the CodeXGLUE CodeCompletion-line dataset (Lu et al., 2021), in addition to the synthetic dataset, SD, used in our initial experiments. Finally, we explore RLHF memorization on different task domains by studying memorization on natural language datasets, LIMA (Zhou et al., 2024) and Anthropic HH (Bai et al., 2022). We summarize our findings below, with full results given in Appendix G:
We observe a slight increase in RL prompt memorization with larger model sizes, with higher rates observed on Gemini Pro across the SD, CodeXGLUE, and LIMA datasets compared to Gemini Nano-1 on SD. However, the memorization rate remains below 1% in all settings. Across all datasets and tasks, a minimal number of reward model training examples exhibit evidence of memorization. This finding suggests that the lack of memorization in reward model training is not specific to the synthetic dataset (SD) or to code completion tasks. Furthermore, memorization rates do not appear to increase substantially with model size. The combination of these results suggest that our findings are robust across different code completion datasets, different task domains, and model scales.
@@ 6 Discussion
As our experimental results have shown, and previous work has supported (Carlini et al., 2022), large language models can memorize training data. However, modern deployed language models go through extensive post-training after pre-training, with most post-training primarily relying on RLHF to align with human preferences. To date, there has been little analysis of the impact of this post-training phase on memorization of training data. Our study performs a careful analysis of the potential for memorization of data introduced in each of the three phases of RLHF. While Rando and Tramèr (2024) examines adversarial poisoning of the RLHF pipeline, our study examines how memorization can arise and propagate among the stages of RLHF in the average case, which is likely more representative of how large model providers currently align their models with user preferences. Additionally, we compare how memorization can occur in algorithms that directly learn from preference data, like IPO (Azar et al., 2024), where no intermediary reward model is trained.
@@
However, we find that algorithms that learn directly from preferences, like IPO, can exhibit much higher memorization than their RLHF counterparts. This is important to note as direct preference learning has gained popularity for its relative simplicity compared to the multi-stage process of RLHF. These results suggest that RLHF is a safer approach than direct preference learning for training on human preference data when concerned about the risk of regurgitation.
It’s important to note that measuring the risk of sensitive data leakage requires more nuance than assessing verbatim memorization because it requires knowledge of exactly what data is considered sensitive. This becomes inherently more difficult with larger, real-world datasets that may contain sensitive information not known ahead of time, in contrast to our work where we create the sensitive data. Additionally, what defines memorization is dependent on the data domain. In the context of code completion, we found that measuring counterfactual memorization was critical, since the highly structured nature of code introduces many potential confounding explanations of why a model produces an output that matches a given target (e.g., the model is a good code completion model). We also found it important to exclude examples where the target was approximately contained in the prompt, since code contains information that is often repeated (e.g. variable names and method calls). Restricting the impact of these confounders on a memorization analysis requires careful consideration of the unique qualities of the data.
@@ A.4 Benchmark evaluation
Additionally, we evaluate our model on the canonical Python coding evaluation of HumanEval introduced by  Chen et al. [2021] that measures correctness for creating programs from docstrings. While not a code completion specific dataset, oftentimes users find value in the ability of a model to generate a function given a function header, docstring, or commented description. This benchmark serves as a proxy for that task. Overall, most models we train are competitive with other small models on code completions tasks (see Appendix I).
@@ Appendix B Instantiation of the working definitions of memorization
Code is highly structured and a model that has learnt language specific style guidelines and has access to relevant contextual information from an examples prompt could conceivably output a completion similar to the target without having memorized the example during training. We control for this by utilizing the counterfactual definition of memorization in Definition 3.1. Comparing completions under the model we are inspecting for memorization and under a control model (that has a separate training dataset) will allow us to rule out matches that are not due to memorization, since the control model has a separate training dataset. We then only count an example as memorized if it is $k$-approximately counterfactually memorized using this fixed control model. Throughout all experiments in Section 5, we use FT.2 as a control model, as this model was not trained on data from SD but performs well on benchmark metrics. We refer to Appendix I for full details of performance and training data for FT.2.
@@ Appendix E Further experiments on RL prompt memorization
Appendix F Results on IPO
These results from Section 5.4 are visualized in Figure 6. In comparison to results on RLHF, preference data memorization is significantly more prevalent with IPO alignment.
(a) Comparison of edit distances between Gemma 2B trained with IPO and the control (Gemma 2B pretrained).
(b) Comparison of edit distances between Gemma 7B trained with IPO and the control (Gemma 7B pretrained).
Figure 6: Normalized edit distance between model completions and targets when models are trained via IPO directly on SD.Base. In Figures 6(a) and 6(b), we compare edit distance between a model trained with IPO directly on SD.Base and its pretrained backbone (that serves as the control as it was not exposed to SD.Base) in its training data). We highlight areas where examples have small edit distances with respect to the control model; these are potential false positives of memorization, and where examples have small edit distances with respect to the model under inspection but not under the control model; these are likely to be true positives of memorization. We find that Gemma-2B and Gemma-7B exhibit 18.2% and 19.5% memorization rates on SD.Base, respectively. Moreover, we see slightly increased memorization at the 7B scale over the 2B scale.
Appendix G Results on different model sizes and tasks
Results across datasets, tasks, model scales (RL prompt data): Figure 7 shows a slight increase in RL prompt memorization with larger model sizes, with higher rates observed on Gemini Pro across the SD, CodeXGLUE, and LIMA datasets compared to Gemini Nano-1 on SD. However, the memorization rate remains below 1% in all settings.
Results across datasets, tasks, model scales (reward model training data): Figure 8 shows that, across all datasets and tasks, a minimal number of reward model training examples exhibit evidence of memorization. This finding suggests that the lack of memorization in reward model training is not specific to the synthetic dataset (SD) or to code completion tasks. Furthermore, memorization rates do not appear to increase substantially with model size.
Figure 7: We measure the memorization rate of RL fine-tuning prompts across different model sizes and datasets. We observe that memorization rate slightly increases when using a larger model, but is still small in absolute terms ($<1\%$ of the dataset).
(a) SD.Base
(b) CodeXGLUE CodeCompletion-line
(c) Anthropic HH.
Figure 8: We use the same set-up as in Figure 3, where we measure the memorization rate of reward model training data from the RL fine-tuned model. We increase the scale of both the reward model (using Gemini Nano instead of T5-Base) and the RL fine-tuned model (using Gemini Pro instead of Gemini Nano), and we use the base Gemini Pro model as the control model. In  Figure 8(a), we use SD.Base as reward model training data, in Figure 8(b), we use the CodeXGLUE CodeCompletion-line dataset as reward model training data (we also use a different split of this dataset for RL fine-tuning), and in  Figure 8(c), we use the Anthropic HH (helpfulness and harmlessness) dataset for reward model training data (and use the LIMA dataset for RL fine-tuning) – both natural language datasets. We subsample the CodeXGLUE CodeCompletion-line dataset and Anthropic HH dataset to be of equal size to SD.Base. We can see that in each experiment, there is relatively little memorization (few points are contained in the blue region).
Appendix H Selection of memorized examples
@@ 12 Returns:
Appendix I Benchmark results and training data provenance summary
@@
Table 1: Training data provenance details for the models we train. All fine-tuned models are initialized from the Gemini Nano-1 (1.8B) pre-trained on 14B tokens from public code repositories as described in Section A.3. SD refers to the synthetic dataset we use for memorization analysis and is described in Appendix J, and PD refers to the public dataset we use to maintain model utility and is described in Section A.1.
@@
Appendix J Construction of the synthetic dataset (SD) and a selection of examples